OpenStudy (anonymous):

On a scale (architects scale) of 1/8" = 1', if a rectangle is drawn 8 1/2" wide x 10 1/4" long, what is the actual length and width?

OpenStudy (anonymous):

8.5 * 8 = 68' wide 10.25 * 8 = 82' in length

OpenStudy (anonymous):

multiple two numbers and the multiply by 64

OpenStudy (anonymous):

the answer is: 68'x82'...but I don't know how to GET TO THIS ANSWER!!! Can anybody help me? Thanks a million!

OpenStudy (anonymous):

Yes, thank you!

OpenStudy (anonymous):

why are you multiplying by 8?

OpenStudy (anonymous):

oh scale is 1/8 means everything is 8 time greater so multiplying by 8 brings the object to scale 8*8 = 64 covers both sides at one time

OpenStudy (anonymous):

Thank you so much! This helps a lot!

OpenStudy (anonymous):

if some thing was drawn 1/4" = 1" then multiply by 4 and so on